OpenStudy (anonymous):

The test scores for a math class are shown below. 81, 85, 82, 93, 85, 84, 95, 87, 88, 91 25. What are the mean, median, and mode of the data set? (1 point)mean: 87; median: 85; mode: 87.3 mean: 85; median: 87.1; mode: 86 mean: 87.3; median: 87; mode; 85 mean: 87.1; median: 86; mode: 85 26. What is the standard deviation of the data set? Round your answer to the nearest tenth. (1 point)2.7 4.4 11.6 19.5 @Luis_Rivera @bnkilinc @kropot72

OpenStudy (kropot72):

The first step is to arrange the data values in ascending order, left to right. Can you please do that an post the result?

OpenStudy (anonymous):

81,82,84,85,85,87,88,91,93,95

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@perl

OpenStudy (kropot72):

Good ordering. To find the mean, add up the 10 values and divide the total by 10. Can you do that?

OpenStudy (kropot72):

@freshkid944 Have you got a result for the mean yet?

OpenStudy (anonymous):

87.1 @kropot72

OpenStudy (anonymous):

thanks!!! now how do I get number 2?

OpenStudy (kropot72):

Good work. The median is the middle value. In this case with 10 values, there is no value in the middle. Therefore we find the average of the two value either side of the center. These value are 85 and 87. Do you know the average of these? It is found by \[\frac{85+87}{2}=?\]

OpenStudy (kropot72):

values*

OpenStudy (anonymous):

128.2 @kropot72

OpenStudy (kropot72):

\[\frac{85+87}{2}=\frac{172}{2}=?\]

OpenStudy (anonymous):

86 @kropot72

OpenStudy (kropot72):

Correct! So we will now find the mode. What value occurs most often in the ordered set of data that you made?

OpenStudy (anonymous):

the mode is 85 :)

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Correct. So we have mean = 87.1, median = 86 and mode = 85. This fits one of the choices. Are you allowed to use a computer or a statistical calculator to find the standard deviation?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Do you have a calculator that can find the standard deviation?

OpenStudy (anonymous):

yes it was 4.4 @kropot72

OpenStudy (kropot72):

Yes, I get 4.4 also.